OpenStudy (anonymous):

solve z=x+iy when sinz=11

OpenStudy (zehanz):

You can try to use the formula:\[\sin z=\frac{ e^{iz}-e^{-iz} }{ 2i }\]So:\[\frac{ e^{iz}-e^{-iz} }{ 2i }=11\]Multiply with 2i and then with \(e^{iz}\) to get a quadratic equation.

OpenStudy (anonymous):

thank you but it says "type your answers in terms of pi and n. type your answers in radians and to 4dp"

OpenStudy (zehanz):

Well, there is still work to do before we get to that: \[e^{iz}-e^{-iz}=22i\]multiply with \(e^{iz}\) and tidy up:\[(e^{iz})^2-22i \cdot e^{iz}-1=0\] Use quadratic formula to find solutions:\[e^{iz}=\frac{ 22i \pm \sqrt{-484+4} }{ 2 }=11i \pm 2 \sqrt{30}i=(11 \pm 2\sqrt{30})i\]

OpenStudy (zehanz):

Now you can solve for z by taking the complex logarithm and dividing by i.

OpenStudy (anonymous):

thank you but i still dont get it what is the answer?

OpenStudy (zehanz):

OK. Definition of complex logarithm: \(\ln z= \ln|z| + i Arg(z)\). Now \(|(11 \pm 2\sqrt{30})i|=11 \pm2\sqrt{30}\) (both numbers are on the positive imaginary axis). Then: Arg(\((11 \pm 2\sqrt{30})i)=\frac{\pi}{2}\). Taking the complex logarithm:\[iz=\ln(11 \pm 2\sqrt{30})+i\frac{\pi}{2}\] Now divide by i:\[z=\frac{\pi}{2}+\frac{\ln(11 \pm 2\sqrt{30})}{i}=\frac{\pi}{2}-i \ln(11 \pm 2\sqrt{30})\]

OpenStudy (zehanz):

We can use a calculator to find decimal values: \(\ln(11+2\sqrt{30})\approx 3.0890\), \(\ln(11-2\sqrt{30})\approx -3.0890\), \(\frac{\pi}{2}\approx 1.5708\) So: \(z \approx 1.5708 \pm 3.0890i\)

OpenStudy (anonymous):

thank you but how is the answer in terms of pi and n?

OpenStudy (zehanz):

We had 1/2 pi as a real part. I'm afraid I don't know what you mean with n.

OpenStudy (anonymous):

im not sure either but thank you

OpenStudy (zehanz):

BTW, it could mean, because sin z is periodic, that the Arg(z) can also be added with n times 2 pi .

OpenStudy (anonymous):

so how would that be included in the answer?

OpenStudy (zehanz):

Hold on, I am looking in to that...

OpenStudy (anonymous):

alright

OpenStudy (zehanz):

It means that \(\frac{\pi}{2}\) must be replaced with \(\frac{\pi}{2}+2n\pi=\frac{(4n+1)\pi}{2}\), so \(z \approx (4n+1)\cdot 1.5708\pm3.0890i\) Hope this helps!

OpenStudy (zehanz):

Checked results on WolframAlpha - worked out fine :) See image.

OpenStudy (anonymous):

thank you